OpenStudy (anonymous):

What is the slope of the line passing through the point (4, −3) and perpendicular to the line y = 2x + 1? Write your answer as a fraction.

OpenStudy (bahrom7893):

slope is -1/2

OpenStudy (anonymous):

That's the answer? that's it?

OpenStudy (bahrom7893):

yea lol if it's perpendicular to y=2x+1, then it's slope is the negative inverse of the slope of y=2x+1, which is 2, so the slope of the new line is -1/2

OpenStudy (anonymous):

OMG thankyou soo much. can you help me with one more?

OpenStudy (bahrom7893):

yessir

OpenStudy (anonymous):

ma'am...

OpenStudy (anonymous):

anyway....Your daily workout plan involves a total of 60 minutes of jogging and biking. You burn 10 calories per minute jogging and 5 calories per minute biking. How many calories will you burn in your 60−minute workout if you jog for 30 minutes?

OpenStudy (bahrom7893):

yes maam

OpenStudy (bahrom7893):

No my daily workout plan does not just involve that.. well actually i'd be dead if it did, it only involves 20 mins of jogging and 40 mins of weightlifting.

OpenStudy (bahrom7893):

anyway...

OpenStudy (bahrom7893):

30 mins jogging, 10 calories per min => 30*10 =300

OpenStudy (anonymous):

lol...wait but i put that and she said it was wrong!

OpenStudy (bahrom7893):

then your question is wrong

OpenStudy (bahrom7893):

or im misunderstanding it

OpenStudy (anonymous):

im gonna put it again...don't worry. how about: Tyrone has a total of 50 DVD's and CD's. If the number of CD's is two more than three times the number of DVD's, how many CD's does he have?

OpenStudy (bahrom7893):

c+d=50 c=3d+2

OpenStudy (anonymous):

So how many CD'S?

OpenStudy (bahrom7893):

sovle for c! come on, it's not hard

OpenStudy (anonymous):

ok hold on

OpenStudy (bahrom7893):

d = 50-c c = 3*d+2

OpenStudy (anonymous):

38?

OpenStudy (bahrom7893):

c = 80 - 3c + 2 4c = 82 c = ..waoah wait..

OpenStudy (anonymous):

what?

OpenStudy (bahrom7893):

i'm getting 20 and a half CDs lol not good...

OpenStudy (anonymous):

lol ikri think its 12

OpenStudy (bahrom7893):

equations are correct.. hold on..

OpenStudy (anonymous):

lol ok

OpenStudy (bahrom7893):

c+d=50 c=3d+2 3d+2+d=50 4d = 48 d = 12 12 DVDS 50 - 12 = 38, YOU WERE RIGHT!! WOOOOHOOO!!!

OpenStudy (anonymous):

YAY!

OpenStudy (anonymous):

An auditorium earned $30,000 in sold−out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

OpenStudy (bahrom7893):

hey im not an answering machine now.... :P

OpenStudy (anonymous):

Just one more i promise!

OpenStudy (bahrom7893):

sunset can finish this one

OpenStudy (anonymous):

u r just hellping every one to day arnt u XD

OpenStudy (anonymous):

whats ur question duell

OpenStudy (anonymous):

An auditorium earned $30,000 in sold−out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

OpenStudy (anonymous):

hi mathy and razor do u want to help

OpenStudy (anonymous):

I need help!

OpenStudy (anonymous):

PLEASE I NEED AN ANSWER I NEED T SUBMIT MY TEST RIGHT NOW

OpenStudy (anonymous):

im sorry its been so long sice iv done this

OpenStudy (anonymous):

ok how about this....A website recorded 1,351 visitors one month. This was 30% less than the number of visitors to the site the previous month. The following equation represents this situation, where x represents the number of visitors to the site the previous month. x − 0.3x = 1,351 How many visitors did the website record in both months combined?

OpenStudy (anonymous):

PLEASE HELLO?

OpenStudy (bahrom7893):

0.7x = 1351 x = 1930

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

how about the before that?